### v1

# Third-order spectrum of gravitational waves sourced by an excited scalar field

Abstract
We calculate for the first time the third-order spectrum of gravitational waves sourced by the amplified scalar field perturbations during inflation using the in-in formalism, and discuss the conditions for the third-order spectrum to be smaller than the second-order one. Assuming an exponential growth of the sub-horizon modes of the scalar field perturbations, we find that the third-order spectrum increases faster than the second-order one as the amplification of the field perturbations increases, and thus the third-order spectrum dominates for detectable gravitational waves, which indicates that the perturbation theory breakdowns.
Introduction. The advanced LIGO detectors opened the era of gravitational-wave (GW) astronomy with the first detection of GWs from a merger of binary black holes Abbott et al. (2016). GW observations allow us to understand the physical processes, such as inflation Boyle and Steinhardt (2008); Guzzetti et al. (2016); Bartolo et al. (2016), happened in the early Universe. Recently, some inflationary models, such as ultra-slow-roll inflation Namjoo et al. (2013); Garcia-Bellido and Ruiz Morales (2017); Germani and Prokopec (2017); Motohashi and Hu (2017); Xu et al. (2020); Fu et al. (2020) and sound speed resonance Cai et al. (2018, 2019, 2020), have been proposed to generate strong GW signals, which are expected to be detected by future laser interferometers.
This Letter focuses on an inflationary scenario in which the scalar field perturbations are amplified during inflation. If the perturbation modes are amplified on small scales due to some mechanisms such as the parametric resonance Cai et al. (2018, 2019, 2020); Zhou et al. (2020); Peng et al. (2021); Cai et al. (2021); Inomata (2022), the modes source GWs through the bilinear source terms during inflation. Meanwhile, small-scale curvature perturbations are also enhanced and then stretched out the horizon during inflaiton. When the modes enter the horizon in the radiation-dominated (RD) era, they induce second-order GWs. Therefore, in such an inflationary scenario there are two peaks in the energy spectrum of GWs Peng et al. (2021).
In the cosmological perturbation theory, tensor perturbations can be expanded as $h_{ij}=h^{(1)}_{ij}+\frac{1}{2}h^{(2)}_{ij}+\frac{1}{6}h^{(3)}_{ij}+\cdot\cdot\cdot$, where $h^{(n)}_{ij}$ is the $n$-th order tensor perturbations. Usually, the two-pint correlation of the second-order tenor perturbations, $\left\langle h^{(2)}_{ij}h^{(2)}_{ij}\right\rangle$, are considered as a good approximation of GWs induced by curvature perturbations in the RD era. Although higher-order contributions are subdominant, they are still observable Yuan et al. (2019); Chang et al. (2023, 2024) under some conditions. For example, the cross-correlation $\left\langle h^{(1)}_{ij}h^{(3)}_{ij}\right\rangle$ can slightly reduce the amplitude of the spectrum of GWs on large scales Chen et al. (2023). Recently, the one-loop power spectrum of GWs generated during inflation is fully considered Ota et al. (2023a, b), where the cross-correlation between the first- and second-order terms contributes to the power spectrum of GWs, especially at the infrared limit. However, the validity of perturbative calculations remains an open question in the cases of ultra-slow-roll Kristiano and Yokoyama (2024a, b) and parameter resonance Inomata (2022).
In this Letter, we calculate for the first time the third-order power spectrum of GWs in the case of the amplified scalar field perturbations during inflation. We only calculate the autocorrelation of third-order terms, i.e., $\left\langle h^{(3)}_{ij}h^{(3)}_{ij}\right\rangle$, corresponding to the third-order power spectrum of GWs. We focus on the infrared (IR) limit of the spectrum because the IR behavior of the spectrum is scale-invariant and is constrained by measurements of the cosmic microwave background. We find that the third-order spectrum increases faster than the second-order one as the amplification of the field perturbations increases and finally dominates, which indicates that the perturbative method has broken down. Moreover, based on the perturbation theory, such a GW signal is undetectable by future GW detectors.
Setup. Some models are proposed to generate large scalar field perturbations, which subsequently sources GWs. However, to provide a model-independent analysis, this paper will not focus on specific models. Instead, we will consider a spectator field $\phi$ minimally coupled to gravity and parameterize the enhancement of its mode functions. We can write down corresponding part of the action as
$S_{\phi}=\frac{M_{\rm p}^{2}}{2}\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{R}+\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{L}_{\phi}.$ (1)
To perturbatively expand this action to higher orders, ADM decomposition of the metric is commonly employed
$\mathrm{d}s^{2}=-N^{2}\mathrm{d}\tau^{2}+\gamma_{ij}\left(\mathrm{d}x^{i}+N^{i}\mathrm{d}\tau\right)\left(\mathrm{d}x^{j}+N^{j}\mathrm{d}\tau\right),$ (2)
where $N$ and $N_{i}$ are the lapse function and shift vector, respectively, serving as Lagrange multipliers, $\tau$ is the conformal time, and $\gamma_{ij}$ is the spatial components of the metric. Since the mapping from background spacetime to perturbed spacetime is not unique, there are gauge freedoms need to be fixed. In our calculations we choose the uniform curvature gauge where the field perturbations $\delta\phi$ act as the dynamical variable. Following Maldacena Maldacena (2003), the spatial parts are written as
$\gamma_{ij}=a^{2}\mathrm{e}^{h_{ij}},$ (3)
where $h_{ij}$ are tensor perturbations which are transverse $\left(\partial^{i}h_{ij}=0\right)$ and traceless $\left(\delta^{ij}h_{ij}=0\right)$. We expand the action to fourth order and solve the constrains to obtain $N$ and $N_{i}$, which are both suppressed by slow-roll parameters in spatially flat gauge Inomata (2022, 2024). As a result, these metric perturbations can be safely neglected. With our definition of $h_{ij}$, we have $\operatorname{det}\left|e^{h_{ij}}\right|=1$, which leads to the fact that tensor perturbations keeps the volume element at non-linear order. Therefore, only kinetic terms of the field Lagrangian contributes to the interaction with tensor parts. Applying Legendre transformation to the action, we arrive at the interaction Hamiltonian
$H_{\text{int }}$ $\equiv H^{(3)}+H^{(4)}$
$=a^{2}\int\mathrm{d}^{3}x\left(-\frac{1}{2}h^{ij}+\frac{1}{4}h^{ik}h_{k}{}^{j}\right)\partial_{i}\delta\phi\partial_{j}\delta\phi.$ (4)
Our goal is calculating the primordial tensor power spectrum
$\left.\left\langle\Omega\left|\sum_{s}h_{\mathbf{q},\rm H}^{s}(\tau)h_{\mathbf{q}^{\prime},\rm H}^{s}(\tau)\right|\Omega\right\rangle\right|_{\tau=0}$
$\equiv(2\uppi)^{3}\delta(\mathbf{q}+{\mathbf{q}^{\prime}})P,$ (5)
where $h_{\mathbf{q},\rm H}^{s}(\tau)$ is a operator in the Heisenberg picture. Vacuum expectation value of the operator $\mathcal{O}$ can be computed with in-in formalism via the following formula
$\langle\mathcal{O}\rangle=\lim_{\tau_{i}\rightarrow-\infty(1-\mathrm{i}\epsilon)}\langle 0|F(\tau;\tau_{i})^{\dagger}\mathcal{O}_{\rm I}(\tau)F\left(\tau;\tau_{i}\right)|0\rangle,$ (6)
where $F(\tau;\tau_{0})$ denotes the evolution operator in the interaction picture, which can be solved iteratively and the Heisenberg picture operator can then be written as
$h_{\mathbf{q},\rm H}^{s}(\tau)=F\left(\tau;\tau_{i}\right)^{\dagger}h_{\mathbf{q}}^{s}F\left(\tau;\tau_{i}\right)$
$=h_{\mathbf{q}}^{s}(\tau)+\mathrm{i}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\left[H_{\mathrm{int},{\rm I}}\left(\tau^{\prime}\right),h_{\mathbf{q}}^{s}(\tau)\right]$
$-\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\int_{\tau_{i}}^{\tau^{\prime}}\mathrm{d}\tau^{\prime\prime}\left[H_{\mathrm{int},{\rm I}}\left(\tau^{\prime\prime}\right),\left[H_{\mathrm{int},{\rm I}}\left(\tau^{\prime}\right),h_{\mathbf{q}}^{s}\right]\right]$
++$\cdots$.+ ⋯ . (7)
With this formula we expand $h_{\mathbf{q},\rm H}^{s}(\tau)$ into many terms consist of the linear order operators. A natural order expansion
$h_{\mathbf{q},\rm H}^{s}(\tau)\equiv h_{\mathbf{q}}^{s,(1)}+h_{\mathbf{q}}^{s,(2)}+h_{\mathbf{q}}^{s,(3)}+\cdots,$ (8)
can be done here with respect to the number of interaction picture operators in each term, for example, $\mathrm{i}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\left[H^{(4)}_{{\rm I}}\left(\tau^{\prime}\right),h_{\mathbf{q}}^{s}(\tau)\right]$ contains in $h_{\mathbf{q}}^{s,(3)}$. Then, using the unitarity of the evolution operator, $F^{\dagger}hhF=(F^{\dagger}hF)(F^{\dagger}hF)$, the expectation value of $hh$ can also be expand as
$\sum_{s}\left\langle\Omega\left|h_{\mathbf{q},\rm H}^{s}(\tau)h_{\mathbf{q}^{\prime},\rm H}^{s}(\tau)\right|\Omega\right\rangle$
$=\lim_{\tau_{i}\rightarrow-\infty(1-\mathrm{i}\epsilon)}\sum_{s}\left(\left\langle h_{\mathbf{q}}^{s,(1)}h_{\mathbf{q}^{\prime}}^{s,(1)}\right\rangle+\left\langle h_{\mathbf{q}}^{s,(2)}h_{\mathbf{q}^{\prime}}^{s,(2)}\right\rangle\right.$
$\left.+\left\langle h_{\mathbf{q}}^{s,(1)}h_{\mathbf{q}^{\prime}}^{s,(3)}+h_{\mathbf{q}}^{s,(3)}h_{\mathbf{q}^{\prime}}^{s,(1)}\right\rangle+\left\langle h_{\mathbf{q}}^{s,(3)}h_{\mathbf{q}^{\prime}}^{s,(3)}\right\rangle+\cdots\right)$
$\equiv\left(P_{11}+P_{22}+P_{13}+P_{33}+\cdots\right)(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{q}^{\prime}\right),$ (9)
[…]
$P_{11}(q)=2\left|v_{q}(0)\right|^{2}=\frac{4H^{2}}{M_{\mathrm{p}}^{2}q^{3}}.$ (10)
In what follows, we consider an exponential growth of the sub-horizon modes of the scalar field perturbations. The amplification of $\delta\phi$ from $t_{i}$ to $t_{f}$ is given by $\delta\phi=\Xi\,\delta\phi^{G}$ with $\Xi=e^{\mu H(t_{f}-t_{i})}$. In terms of the conformal time, the amplification factor becomes $\Xi=(\tau_{i}/\tau_{f})^{\mu}$, where $\tau_{i}$ and $\tau_{f}$ correspond to $t_{i}$ and $t_{f}$, respectively. In the momentum space the mode functions are given by $u_{p}(\tau)=\Xi(\tau)u_{p}^{G}(\tau)$ with
$\Xi(\tau)=\left\{\begin{aligned} &1&\tau<\tau_{i},\\ &\left(\frac{\tau_{i}}{\tau}\right)^{\mu}&\tau_{i}\leq\tau\leq\tau_{f},\\ &\left(\frac{\tau_{i}}{\tau_{f}}\right)^{\mu}&\tau>\tau_{f}.\end{aligned}\right.$ (11)
Second-order spectrum. We present a quick review of the one-loop calculations with the commutator language. We introduce two Green’s functions $G^{G}_{q}$ and $G_{q}$ to simplify the expressions later in this paper
$\left[h_{\mathbf{p}}^{s}\left(\tau^{\prime}\right),h_{\mathbf{q}}^{s_{1}}(\tau)\right]$
$=(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{p}\right)\left(v_{q}\left(\tau^{\prime}\right)v_{q}^{*}(\tau)-v_{q}^{*}\left(\tau^{\prime}\right)v_{q}(\tau)\right)\delta^{ss_{1}}$
$\equiv(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{p}\right)\frac{4}{M_{\rm p}^{2}}\frac{\mathrm{i}}{a(\tau^{\prime})^{2}}G_{q}^{G}\left(\tau;\tau^{\prime}\right)\delta^{ss_{1}},$ (12)
and
$\left[\delta\phi_{\mathbf{p}}\left(\tau^{\prime}\right),\delta\phi_{\mathbf{q}}(\tau)\right]$
$=(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{p}\right)\left(u_{q}\left(\tau^{\prime}\right)u_{q}^{*}(\tau)-u_{q}^{*}\left(\tau^{\prime}\right)u_{q}(\tau)\right)$
$\equiv(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{p}\right)\frac{\mathrm{i}}{a(\tau^{\prime})^{2}}G_{q}\left(\tau;\tau^{\prime}\right),$ (13)
Working out all those commutators, we arrive at
$h_{\mathbf{q}}^{s,(2)}=\frac{2}{M_{\rm p}^{2}}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}G_{q}^{G}\left(\tau;\tau^{\prime}\right)S_{\mathbf{q}}^{s}\left[\delta\phi\left(\tau^{\prime}\right)\right],$ (14)
where $S_{\mathbf{q}}^{s}$ represents the source term and it is defined by
$S_{\mathbf{q}}^{s}\left[\delta\phi\left(\tau^{\prime}\right)\right]\equiv\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}e_{ij}^{s*}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}\delta\phi_{\mathbf{p}}\delta\phi_{\mathbf{q}-\mathbf{p}},$ (15)
where $e_{ij}^{s*}(\hat{q})$ is the polarization tensor. From $h_{\mathbf{q}}^{s,(2)}$, one can obtain the $P_{22}$ which is exactly the power spectrum of the so called scalar-induced GWs during inflation. Meanwhile, As emphasized in Ota et al. (2023a, b), $P_{13}$ cannot be neglected since it is also of one-loop order, just as $P_{22}$ does. To compute this cross-correlation term, $h_{\mathbf{q}}^{s,(3)}$ must be considered, and the results are given by
$h_{\mathbf{q}}^{s,(3)}=h_{\mathbf{q}}^{s,(3a)}+h_{\mathbf{q}}^{s,(3b)},$ (16a)
$h_{\mathbf{q}}^{s,(3a)}=\frac{2}{M_{\rm p}^{2}}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\int_{\tau_{i}}^{\tau^{\prime}}\mathrm{d}\tau^{\prime\prime}G_{q}^{G}\left(\tau;\tau^{\prime}\right)$
×$\times\int\frac{\mathrm{d}^{3}p\mathrm{d}^{3}k}{(2\uppi)^{6}}e_{ij}^{s*}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}G_{p}\left(\tau^{\prime};\tau^{\prime\prime}\right)\sum_{s_{1}}h_{\mathbf{k}}^{s_{1}}(\tau^{\prime\prime})$
×$\times e_{mn}^{s_{1}}(\hat{k})\mathbf{p}_{m}\mathbf{p}_{n}\{\delta\phi_{\mathbf{p}-\mathbf{k}}(\tau^{\prime\prime})\delta\phi_{\mathbf{q}-\mathbf{p}}(\tau^{\prime})\},$ (16b)
$h_{\mathbf{q}}^{s,(3b)}=-\frac{2}{M_{\rm p}^{2}}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}G_{q}^{G}\left(\tau;\tau^{\prime}\right)\int\frac{\mathrm{d}^{3}p\mathrm{d}^{3}k}{(2\uppi)^{6}}\sum_{s_{2}}e_{ik}^{s*}(\hat{q})$
×$\times e_{kj}^{s_{2}}(\hat{p})\mathbf{k}_{i}(\mathbf{q}-\mathbf{k})_{j}\delta\phi_{\mathbf{k}}(\tau^{\prime})\delta\phi_{\mathbf{q}-\mathbf{p}-\mathbf{k}}(\tau^{\prime})h_{\mathbf{p}}^{s_{2}}(\tau^{\prime}),$ (16c)
where the expression is divided into two parts corresponding to $H^{(3)}$ and $H^{(4)}$, respectively. Using these results, we obtain integral expressions for the one-loop power spectrum through several contractions. We first address the momentum integrals, which involve a triple integral over the entire momentum space.
Following Ota et al. (2023b), we expect that the dominant contributions arise from the enhanced modes, as other modes are in ground state. Therefore, contributions from those ground state modes can be approximately neglected thus we only integrate over momentum interval $p_{*}$ to $p_{*}+d$. By applying appropriate approximations to the integration region, we analytically perform these momentum integrals. The details of our approximations are provided in the Appendix A. Finishing momentum integrals yields the integrand of the time integral, which we can evaluate numerically. We fix the initial and final time of the growth while varying the index $\mu$ to adjust the magnification. In order to figure out the momentum dependence of one-loop order power spectrum, we have chosen a specific magnification index $\mu$ and calculated $P_{13(a(b))}$, which corresponds to correlations between $h_{\mathbf{q}}^{s,1}$ and $h_{\mathbf{q}}^{s,(3a(b))}$ respectively as shown in Fig. 1.
Figure 1: One loop order tensor power spectrum including $P_{22}$ and $P_{13}$, where they have been normalized by dividing $P_{11}$. In this figure, we denote positive values with solid lines and negative values with dashed lines. Initial and final times of the growth are fixed to $x_{i}=p_{*}\tau_{i}=-30$ and $x_{f}=p_{*}\tau_{f}=-3$, respectively, while we take $\mu=3.2$. Also, we also used the normalized momentum $\tilde{q}=\frac{q}{p_{*}}$.
As illustrated by Ota et al. (2023b), there are several noteworthy aspects about these results. We are already familiar with the $\mathcal{O}(q^{3})$ infrared (IR) behavior of the $P_{22}$ terms, which arises naturally from the causal production of GWs. However, the $P_{13}$ terms exhibit scale invariance for IR modes. To explore the IR behavior, we consider the limit where $|q\tau|\ll 1$. In this regime, Green function can be approximated as
$G^{G}_{q}\left(0;\tau^{\prime}\right)=-\frac{4H^{2}a^{\prime 2}\tau^{\prime 3}}{3M_{\mathrm{pl}}^{2}}\left[1+\mathcal{O}\left(\tau^{\prime 2}q^{2}\right)\right].$ (17)
Further analysis reveals that for $q\ll p_{*}$ and $|q\tau|\ll 1$, $|q\tau^{\prime}|\ll 1$, the expression $S_{\mathbf{q}}^{s}[\delta\phi]$ can be written as $e^{ij,s,*}(\hat{q})\Sigma_{ij}[\delta\phi]$, where $\Sigma_{ij}[\delta\phi]$ is almost scale-invariant. Consequently, we find that $\frac{P_{22}}{P_{11}}=\mathcal{O}\left(\tilde{q}^{3}\right)$ on large scales. Notably, there is an “additional” term of the form $v_{\mathbf{q}}(\tau)v^{*}_{\mathbf{q}}(\tau^{\prime})$ in the expression for $P_{13}$, which exactly cancels out the $q$ dependence introduced when dividing by $P_{11}$, resulting in a $\mathcal{O}\left(\tilde{q}^{0}\right)$ IR behavior.
It is noteworthy that $P_{13}\approx P_{13(a)}$, particularly in the IR limit. This observation indicates that the traditional reliance on the number of loops for quantifying orders of magnitude may be misleading. A more precise approach involves considering the exponent of the ratio ${H}/{M_{\rm p}}$ alongside the enhancement factor $\Xi$. For instance, $P_{13(a)}$ is of order $\left({H}/{M_{\rm p}}\right)^{2}\Xi^{4}$ which is of same order as $P_{22}$, we can mark their order by $(2,4)$, while $P_{13(b)}$ is only of order $(2,2)$. For the large-scale spectrum, it is straightforward to verify that this approximation aligns well with the numerical results, effectively capturing the underlying behavior.
Third-order spectrum. The counter-correlation terms exhibit an interesting characteristic: Fig. 1 clearly demonstrates that the large-scale values of $P_{13}$ is significantly larger, even exceeding the peak value of $P_{22}$, which contrasts sharply with the corresponding terms in the RD era calculated in Chen et al. (2023). The significant magnitude of $\left|\left\langle h_{\mathbf{k}}^{(1)}h_{\mathbf{-k}}^{(3)}\right\rangle\right|$ indicates that maybe the $\left\langle h_{\mathbf{k}}^{(3)}h_{\mathbf{-k}}^{(3)}\right\rangle$ terms themselves can not be safely neglected. These terms consist of three types: $P_{33aa}$, $P_{33ab}$, and $P_{33bb}$, which are of order $(4,8)$, $(4,6)$, and $(4,4)$, respectively, suggesting that we can have following approximation $P_{33}\approx P_{33aa}$. Upon meticulously handling all these contractions, we identify three types of momentum structures, which can be represented diagrammatically in Fig. 2.
Figure 2: Three diagrams correspond to $P_{33aa}$
It is easy to identify that only Fig. 2 has a $\mathcal{O}\left(\tilde{q}^{0}\right)$ momentum dependence. The reason is that the momentum of the inner tensor propagator is fixed to $q$ by the conservation of 3-momentum, which leaves a $v_{q}(\tau^{\prime})v^{*}_{q}(\tau^{\prime\prime})$ term in the expressions. For the other two diagrams, the momentum of inner $h$ is integrated over the entire momentum space so they will obtain a $\tilde{q}^{3}$ depression for large scales after being normalized by $P_{11}$.
Now let us consider the third order contribution to the GW power spectrum from the Fig. 2. The expression can be written as
$\left\langle h_{\mathbf{q}}^{r}h_{\mathbf{q}^{\prime}}^{s}\right\rangle=(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{q}^{\prime}\right)\frac{4}{M_{\rm p}^{4}}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\int_{\tau_{i}}^{\tau^{\prime}}\mathrm{d}\tau^{\prime\prime}$
×$\times G_{q}^{G}\left(\tau;\tau^{\prime}\right)\int_{\tau_{i}}^{\tau_{1}}\mathrm{d}\tau_{1}^{\prime}\int_{\tau_{i}}^{\tau_{1}^{\prime}}\mathrm{d}{\tau_{1}}^{\prime\prime}G_{q}^{G}\left(\tau_{1};\tau_{1}^{\prime}\right)v_{q}(\tau^{\prime\prime})v^{*}_{q}(\tau_{1}^{\prime\prime})$
×$\times\sum_{s_{1}}\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}2e_{ij}^{r}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}e_{mn}^{s_{1}^{*}}(\hat{q})\mathbf{p}_{m}\mathbf{p}_{n}G_{q}\left(\tau^{\prime};\tau^{\prime\prime}\right)$
×$\times\Re(u_{|\mathbf{q}-\mathbf{p}|}(\tau_{1}^{\prime\prime})u_{|\mathbf{q}-\mathbf{p}|}^{*}(\tau_{1}^{\prime}))\int\frac{\mathrm{d}^{3}k}{(2\uppi)^{3}}e_{\alpha\beta}^{s*}(\hat{q})\mathbf{k}_{\alpha}\mathbf{k}_{\beta}e_{kl}^{s_{1}}(\hat{q})$
×$\times\mathbf{k}_{k}\mathbf{k}_{l}G_{q}\left(\tau_{1}^{\prime};\tau_{1}^{\prime\prime}\right)2\Re\left[u_{|\mathbf{q}-\mathbf{k}|}(\tau_{1}^{\prime\prime})u_{|\mathbf{q}-\mathbf{k}|}^{*}(\tau_{1}^{\prime})\right],$ (18)
Following the method used to calculate the one-loop contributions, the results are straightforward.
Figure 3: Infrared values of $P_{13}$ and $P_{33}$, where the spectrum have also been normalized.
In Fig. 3, we compare the IR values of $P_{13}$ and $P_{33}$ for various $\mu$, where we can clearly find that when $P_{13}$ are strictly smaller than linear terms, two-loop contributions are even much smaller. However, as we gradually enhance the scalar perturbations, $P_{33}$ is more intensely enhanced and finally exceed $P_{13}$. There is no doubt that this result agree well with the fact that, as the scalar modes become larger $P_{33}$ grows as $\Xi^{4}$ order and $P_{13}$ grows as $\Xi^{2}$ order.
Conclusion and discussion. As we have mentioned above, the IR value of $P_{13}$ is always larger than $P_{22}$, so if $\delta\phi$ are enhanced large enough to make the third order corrections even larger, we have to admit that $h_{\mathbf{q}}^{s,(3)}$ can’t be neglected in those circumstances, which means $P_{22}$ can not be consider as a good approximation of the physical observable. More numerical experiments also showed that $P_{22}$ can not even exceed the linear order spectrum with the request that the GWs are perturbatively induced. While we have proved that the third order terms are larger during inflation through direct calculations, one may wonder why this did not happen in the RD era. The key point is that during inflation, the Green function of scalar modes are of $\Xi^{2}$ order, while when we consider the scalar-induced GWs in the RD era, only the initial condition not the Green functions are enhanced.
In more realistic models, this kind of enhancement corresponds to changing of the linear-order equations of motions (EoMs), for example, in the widely considered resonance models Cai et al. (2018, 2019, 2020); Zhou et al. (2020); Peng et al. (2021); Inomata (2022), a tiny oscillating structure is added to the inflaton potential which contributes a resonance term in the EoM of inflaton perturbations. On the other hand, in our parameterization, during the period of time that the oscillating term is active, the mode function of the scalar perturbations will be exponentially enhanced, which is similar to those resonance models. However, we are not sure whether the results will change for more realistic models, we will do further theoretical and numerical examinations in our future works.
Actually, there is another perspective of our results, if we consider the higher order corrections with the EoM method, corresponding three order tensor modes can be expressed as
${h_{\mathbf{q}}^{s}}^{(3)}=\frac{2}{M_{\rm p}^{2}}\int^{\tau}\mathrm{d}\tau^{\prime}G^{\rm BD}_{k}(\tau,\tau^{\prime})$
×$\times\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}e_{ij}^{s*}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}\delta\phi_{\mathbf{p}}^{(2)}\delta\phi_{\mathbf{q}-\mathbf{p}},$ (19)
where
$\delta\phi_{\mathbf{q}}^{(2)}=\int^{\tau}\mathrm{d}\tau^{\prime}G_{k}(\tau,\tau^{\prime})$
×$\times\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}\sum_{s}e_{ij}^{s}(\hat{p})\mathbf{q}_{i}\mathbf{q}_{j}h_{\mathbf{p}}^{s}\delta\phi_{\mathbf{q}-\mathbf{p}}.$ (20)
We can surprisingly find that the order of $\left\langle\delta\phi_{-\mathbf{p}}^{(2)}\delta\phi_{\mathbf{p}}^{(2)}\right\rangle$ is the same as $\left\langle h_{-\mathbf{p}}^{(2)}h_{\mathbf{p}}^{(2)}\right\rangle$, so if the power spectrum of the induced GWs are much larger than $P_{11}$, the scalar perturbation itself is also non-perturbative. It is also instructive here to compare our calculations with the results presented in Inomata et al. (2023), where the self-interaction of the scalar field perturbations are considered. These interactions also contributes considerable loop corrections when the scalar modes are enhanced large enough and thus give perturbative constraints on the scalar magnification. Although both of them considered higher-order scalar two-point correlation functions, we argue that the effect we are discussing is more general because it is an intrinsic property of the GW production, while the calculations of Inomata et al. (2023) rely on the details of the resonance models. In this letter, we proved directly that non-perturbative effects come from exactly the same mechanism. It is also worth noticing that recent studies on small-scale nonlinear processes during inflation have advanced to the level of lattice simulations Caravano et al. (2024), however, future works are still needed to perform lattice simulations of the tensor modes.
More Intuitively, the Fig. 2 of $\left\langle h_{-{\mathbf{p}}}^{(3)}h_{\mathbf{p}}^{(3)}\right\rangle$ is directly related to the self-correlations of the second order $\phi$; its values are shown below, where we can also see the two-loop spectrum exceed the one-loop spectrum as the scalar perturbations become larger.
Figure 4: Power spectrum corresponds to Fig. 2 of $P_{33}$ and $P_{22}$, where the spectrum and momentum are also normalized.
It should be mentioned here that we haven’t considered all two-loop diagrams, which also include cross-correlation terms like $\left\langle h_{-\mathbf{p}}^{(2)}h_{\mathbf{p}}^{(4)}\right\rangle$ and $\left\langle h_{-\mathbf{p}}^{(1)}h_{\mathbf{p}}^{(5)}\right\rangle$. However, using the order parameter we mentioned above, we find that they are of the same order, besides, for higher order diagrams we found that they contain more enhancement factor so they may grow even faster than two loop terms. These facts indicate that perturbative calculations for observable scalar induced GWs during inflation era may be impossible.
Acknowledgements. We thank Xingang Chen, Keisuke Inomata, Jason Kristiano and Atsuhisa Ota for useful comments and discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201501, in part by the National Natural Science Foundation of China under Grant No. 12475067 and No. 12235019.
Appendix A Analytical calculation of momentum integral in one-loop contribution
A.1 Linear order
The Fourier modes of the scalar field perturbations and tensor perturbations are expressed as
$\delta\phi(\tau,\mathbf{x})$ $=\int\frac{\mathrm{d}^{3}k}{(2\uppi)^{3}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{x}}\delta\phi_{\mathbf{k}}(\tau)$ (21)
$h_{ij}(\tau,\mathbf{x})$ $=\int\frac{\mathrm{d}^{3}k}{(2\uppi)^{3}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{x}}\sum_{s=\pm 2}e_{ij}^{s}(\mathbf{k})h_{\mathbf{k}}^{s}(\tau).$ (22)
Taken only quadratic Hamiltonian in to consideration, we can first solve the time evolution of $\phi_{\mathbf{k}}$ and $h_{\mathbf{k}}^{s}$ in the interaction picture. According to standard quantization procedures, these perturbation terms can be expand in to creation and annihilation operators as
$\delta\phi_{\mathbf{q}}(\tau)$ $=u_{q}(\tau)a_{\mathbf{q}}+u_{q}^{*}(\tau)a_{-\mathbf{q}}^{\dagger},$ (23)
$h_{\mathbf{q}}^{s}(\tau)$ $=v_{q}(\tau)b_{\mathbf{q}}^{s}+v_{q}^{*}(\tau)b_{-\mathbf{q}}^{s\dagger},$ (24)
where the $u_{q}(\tau)$ and $v_{q}(\tau)$ are mode functions, and the annihilation and creation operators satisfy
$a_{\mathbf{q}}|0\rangle=b_{\mathbf{q}}^{s}|0\rangle=0$ (25)
${\left[a_{\mathbf{q}},a_{-{\mathbf{q}^{\prime}}}^{\dagger}\right]}=(2\uppi)^{3}\delta(\mathbf{q}+{\mathbf{q}^{\prime}})$ (26)
${\left[b_{\mathbf{q}}^{s},b_{-{\mathbf{q}^{\prime}}}^{r\dagger}\right]}=(2\uppi)^{3}\delta^{sr}\delta(\mathbf{q}+{\mathbf{q}^{\prime}}).$ (27)
For ground states that the scalar perturbations are not amplified, the mode functions are given by
$u_{k}^{\mathrm{G}}(\tau)$ $=\frac{H}{\sqrt{2k^{3}}}(1+\mathrm{i}k\tau)\mathrm{e}^{-\mathrm{i}k\tau},$ (28)
$v_{k}^{\mathrm{G}}(\tau)$ $=\frac{2H}{M_{\mathrm{p}}\sqrt{2k^{3}}}(1+\mathrm{i}k\tau)\mathrm{e}^{-\mathrm{i}k\tau}.$ (29)
Hence we can get the linear order power spectrum through direct calculations
$P_{11}(q)=2\left|v_{q}(0)\right|^{2}=\frac{4H^{2}}{M_{\mathrm{p}}^{2}q^{3}}.$ (30)
A.2 Interaction picture
The evolution operator in the interaction picture are defined as
$F\left(\tau;\tau_{i}\right)$ $=U_{0}^{-1}\left(\tau,\tau_{0}\right)U\left(\tau,\tau_{0}\right)$ (31)
where the $U_{0}\left(\tau,\tau_{0}\right)$ and $U\left(\tau,\tau_{0}\right)$ denote evolution operator correspond to free and full hamiltonian respectively. We can get directly from definition that this operator satisfy
$\partial_{\tau}\left[U_{0}^{-1}\left(\tau,\tau_{0}\right)U\left(\tau,\tau_{0}\right)\right]=-\mathrm{i}U_{0}^{-1}\left(\tau,\tau_{0}\right)U\left(\tau,\tau_{0}\right)H_{\mathrm{int}},$ (32)
which can be solved iteratively
$F\left(\tau;\tau_{i}\right)$ $=\mathbb{T}\exp\left(-\mathrm{i}\int_{\tau_{0}}^{\tau}\mathrm{d}\tau^{\prime\prime}H_{\mathrm{int},{\rm I}}\left(\tau^{\prime\prime}\right)\right)$
$=1-\mathrm{i}\lambda\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}H_{\mathrm{int},{\rm I}}\left(\tau^{\prime}\right)$
$-\lambda^{2}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}\int_{\tau_{i}}^{\tau^{\prime}}\mathrm{d}\tau^{\prime\prime}H_{\mathrm{int},{\rm I}}\left(\tau^{\prime}\right)H_{\mathrm{int},{\rm I}}\left(\tau^{\prime\prime}\right)$
$+\mathcal{O}\left(\lambda^{3}\right),$ (33)
note that the order counting parameter $\lambda$ is shown explicitly.
A.3 Momentum integrals
As mentioned in the previous chapter, only enhanced k-modes are included in our momentum integrals. Here, I would like to used two examples to illustrate our method. First, the momentum part of $P_{22}$ can be written as
$\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}e^{s*}(q,p)e^{s^{\prime}}(q,p)u_{p}(\tau^{\prime})f(|\mathbf{q}-\mathbf{p}|,p),$ (34)
where $e^{s*}(q,p)=e_{ij}^{s*}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}$ and
$f(|\mathbf{q}-\mathbf{p}|,p)=u_{{p}}^{*}(\tau^{\prime\prime})u_{|\mathbf{q}-\mathbf{p}|}(\tau^{\prime})u_{|\mathbf{q}-\mathbf{p}|}^{*}(\tau^{\prime\prime}).$ (35)
At first glance, a reasonable approximation for a narrow spectrum is
$f\left(|\mathbf{q}-\mathbf{p}|,p\right)=\delta\left[\ln\left(\frac{|\mathbf{q}-\mathbf{p}|}{p_{*}}\right)\right]\delta\left[\ln\left(\frac{p}{p_{*}}\right)\right]f\left(p_{*},p_{*}\right).$ (36)
With this condition, the momentum integral can be simplified as a angle integral
$\int\frac{\mathrm{d}\theta}{(2\uppi)^{3}}p_{*}^{8}\uppi\frac{\sin^{4}{\theta}}{q}\delta\left(\theta-\theta_{0}\right)=\frac{p_{*}^{4}\uppi}{q}\left(p_{*}^{2}-\frac{q^{2}}{4}\right)^{2}.$ (37)
This result will finally leads us to a $\mathcal{O}(q^{2})$ infrared behaviour which should not happen in local production of GWs. We consider a more physical setting to reveal the problems hidden behind the $\delta$ approximation and to obtain a more reasonable result. We require that only from a specific $p$-mode $p_{*}$ to $p_{*}+d$, the mode functions of $\phi$ are enhanced, thus we have two characteristic momentum scale $q$ and $d$. We consider the integration region $\Omega$ in the momentum space, it is the overlap part of two spherical shell as shown in Fig. 4.
$\Delta\Pi=\int_{\Omega}\frac{\mathrm{d}p\mathrm{d}\theta}{(2\uppi)^{3}}\uppi p^{6}\sin^{5}{\theta}$ (38)
For $d\ll q$, the integration region is a ring and the polar angle only vary in a very narrow interval around $\theta_{0}$, which is similar to the $\delta$ case. However, no matter how small the width $d$ is, there always exist $q<d$, and for these infrared $q$ we must integrate over a shell which means $\theta$ vary from 00 to $\uppi$. It still difficult to obtain the exact value of this integral, so we decided to approximate it from two limits and link them up at $q=d$, finally we adopt the following results
$\Delta\Pi=\left\{\begin{aligned} &\frac{\uppi p_{*}^{6}d^{2}\sin^{4}{\theta_{0}}}{(2\uppi)^{3}q},&\left(q<d\right)\\ &\frac{\uppi p_{*}^{6}d}{(2\uppi)^{3}},&\left(q>d\right).\end{aligned}\right.$ (39)
and now the momentum integral of $P_{22}$ have been completed. In the second example we are going to deal with the momentum part of Fig. 2 of $P_{33aa}$. We first give the full expressions of $\left\langle h_{-\mathbf{p}}^{(3)}h_{\mathbf{p}}^{(3)}\right\rangle^{(b)}$
$\left\langle h_{\mathbf{q}}^{r}h_{\mathbf{q}^{\prime}}^{s}\right\rangle=(2\uppi)^{3}\delta\left(\mathbf{q}+\mathbf{q}^{\prime}\right)\frac{4}{M_{\rm p}^{4}}\int_{\tau_{i}}^{\tau}\mathrm{d}\tau^{\prime}G_{q}^{G}\left(\tau;\tau^{\prime}\right)$
$\int_{\tau_{i}}^{\tau}\mathrm{d}\tau_{1}^{\prime}G_{q}^{G}\left(\tau;\tau_{1}^{\prime}\right)\int\frac{\mathrm{d}^{3}p}{(2\uppi)^{3}}e_{ij}^{r}(\hat{q})\mathbf{p}_{i}\mathbf{p}_{j}e_{mn}^{s^{*}}(\hat{q})\mathbf{p}_{m}\mathbf{p}_{n}$
$u_{|\mathbf{q}-\mathbf{p}|}(\tau^{\prime})u_{|\mathbf{q}-\mathbf{p}|}^{*}(\tau_{1}^{\prime})U(\tau^{\prime},\tau_{1}^{\prime}),$ (40)
where
$U(\tau^{\prime},\tau_{1}^{\prime})=4\int_{\tau_{i}}^{\tau^{\prime}}\mathrm{d}\tau^{\prime\prime}\int_{\tau_{i}}^{\tau_{1}^{\prime}}\mathrm{d}{\tau_{1}}^{\prime\prime}G_{p}\left(\tau^{\prime};\tau^{\prime\prime}\right)G_{p}\left(\tau_{1}^{\prime};\tau_{1}^{\prime\prime}\right)$
×$\times\int\frac{\mathrm{d}^{3}k}{(2\uppi)^{3}}\sum_{s_{2}}|e_{\alpha\beta}^{s_{2}}(\hat{k})\mathbf{p}_{\alpha}\mathbf{p}_{\beta}|^{2}v_{k}(\tau^{\prime\prime})$
×$\times v^{*}_{k}(\tau_{1}^{\prime\prime})u_{|\mathbf{p}-\mathbf{k}|}(\tau^{\prime\prime})u_{|\mathbf{p}-\mathbf{k}|}^{*}(\tau_{1}^{\prime\prime}).$ (41)
Noticed that the integral of $p$ is similar to the one-loop diagram, we will focus on the $k$ integral which is independent of $q$ so the $\delta$ approximation can be safely applied, but this time the normalization factor have been chosen to be $d$. After some direct calculations, only a integral of $\theta$ is left
$Y=\int_{\frac{\uppi}{2}}^{\uppi}\mathrm{d}\theta\frac{\sin^{5}{\theta}}{\cos^{2}{\theta}}(1+\mathrm{i}k\tau^{\prime\prime})(1-\mathrm{i}k\tau_{1}^{\prime\prime})\mathrm{e}^{\mathrm{i}k(\tau_{1}^{\prime\prime}-\tau^{\prime\prime})},$ (42)
where $k=-2p_{*}\cos{\theta}$. Unfortunately $Y$ goes to infinity as $\theta$ approaches ${\uppi}/{2}$, which means we are facing a infrared divergence and we have to regularize it. To do that we introduce a positive parameter $\theta_{0}$ by hand and integrate it analytically
$Y(\theta_{0})$ $=\int_{\frac{\uppi}{2}+\theta_{0}}^{\uppi}\mathrm{d}\theta\frac{\sin^{5}{\theta}}{\cos^{2}{\theta}}(1+\mathrm{i}k\tau^{\prime\prime})(1-\mathrm{i}k\tau_{1}^{\prime\prime})\mathrm{e}^{\mathrm{i}k(\tau_{1}^{\prime\prime}-\tau^{\prime\prime})}$
$=-\frac{1}{\theta_{0}}+Y_{0}+Y_{1}\theta_{0}+\mathcal{O}(\theta_{0}^{2}).$ (43)
To do complete the regularization process we just need to subtract the divergent term and take $\theta_{0}\rightarrow 0$ so that $Y_{0}$ is just what we need.